4x+7y=-5xy solve for y: 4x+7y=-5xy solve for y
Answer by orca(336)

4x+7y=-5xy
4x+7y-7y=-5xy-7y (subtracting 7y from both sides)
4x=-5xy-7y
4x=y(-5x-7)
4x/(-5x-7)=y (dividing both sides by (-5x-7))
So y =4x/(-5x-7)
Or written as:
y = 4x/(-(5x+7))=-4x/(5x+7)